If the circumradius of an isoceless triangle ABC is equal to AB (= AC), then angle A is equal to

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
sin B = b / 2R
= AC / 2
= R / 2R [Given AB = AC = R]
= 1/2

B = p / 6 or 5p / 6
When B = 5p / 6, C = 5p / 6 [AB = AC ? B = C]
?B + C > p
So, B = 5p / 6 not possible
Hence, B should bep / 6,C = p / 6
A = p - (p / 6 + p / 6)
A = 2p / 3
Thanks
Bharat Bajaj
askiitians faculty
IIT Delhi